# On the Price of Decentralization in Decentralized Detection

### v1

@@ A-E Proof of Theorem 5
For the product distribution of the nodes, let $x=(x^{(1)}_{1},x{(t)}_{2},\dots,x^{(t)}_{n})\in\mathcal{X}^{t}$, and we have
@@
$\beta^{(t)}_{i}(\xi;\eta)=$ $\;P_{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(X^{(t)}_{j})}{P_{j,0}(X^{(t)}_{j})}\leq\eta\right\}$
@@
(28)$=\sum_{x\in\mathcal{X}^{t}}$ $\;P^{\otimes t}_{\theta^{*}}(x)\exp\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,\theta}(x^{(t)}_{j})}\right\}\mathbbm{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}\leq\eta\right\}.$ (29)
@@
$\sum_{j=1}^{n}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,\theta}(x^{(t)}_{j})}=(1-\theta^{*})\left(\sum_{j=1}^{n}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}\right)+\mathrm{CI}(P_{0},P_{1}).$
@@
(29)$=\sum_{x\in\mathcal{X}^{t}}$ $\;P^{\otimes t}_{\theta^{*}}(x)\exp\left\{(1-\theta^{*})\left(\sum_{\tau=1}^{t}\sum_{j=1}^{n}\log\frac{P_{j,0}(x^{(t)}_{j})}{P_{j,\theta}(x^{(t)}_{j})}\right)-\mathrm{CI}t\right\}\mathbbm{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}\leq\eta\right\}.$ (30)
Similarly to the trick in Section A-D, we have
@@
$=$ $\;\sum_{\tau=1}^{t}\sum_{j=1}^{n}(1-\theta^{*})\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}+\sum_{\tau=1}^{t}\sum_{j=1}^{n}\left(1-(1-\theta^{*})\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}$
$\leq$ $\;\sum_{\tau=1}^{t}\sum_{j=1}^{n}(1-\theta^{*})\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}+\underbrace{(1-\theta^{*})\frac{\rho}{1-\rho}\sqrt{\frac{1-\pi_{i}}{\pi_{i}}\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}L^{2}_{j}\right)}}_{C_{i}}.$
@@
By Lemma 3, we know that $Y^{(t)}_{i}$’s CDF converges to the one of the standard normal distribution with some remaining terms, and so far we have equation 31 become
$\;\exp\left\{-\mathrm{CI}t+(1-\theta^{*})C_{i}\right\}\sum_{y\in\mathcal{Y}}P_{Y^{(t)}_{i}}(y)\bigg{\{}\exp\left\{(1-\theta^{*})S^{(t)}_{n}\sqrt{t}y\right\}\mathbbm{1}\left\{(1-\theta^{*})S^{(t)}_{n}\sqrt{t}yx\leq(1-\theta^{*})\eta\right\}\bigg{\}}$
$=$ $\;\exp\left\{-\mathrm{CI}t+(1-\theta^{*})\eta+(1-\theta^{*})C_{i}\right\}\sum_{z\leq 0}P_{Y^{(t)}_{i}}\left(\frac{z+(1-\theta^{*})\eta}{(1-\theta^{*})S^{(t)}_{n}\sqrt{t}}\right)e^{z}.$
@@
using the similar method in Section A-D. Thus, the type-II error probability is
$\beta^{(t)}_{i}(\xi;\eta)=\exp\left\{-\mathrm{CI}t-\frac{1}{2}\log t+(1-\theta^{*})\eta+(1-\theta^{*})C_{i}-\log\left((1-\theta^{*})\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\right\}.$ (32)
@@
$\alpha^{(t)}_{i}(\xi;\eta)=\exp\left\{-\mathrm{CI}t-\frac{1}{2}\log t-\theta^{*}\eta+\theta^{*}C_{i}-\log\left(\theta^{*}\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\right\}.$ (33)
@@
$\bar{\alpha}^{(t)}_{i}(\xi;\eta)=\exp\left\{-\mathrm{CI}t-\frac{1}{2}\log t-\theta^{*}\eta+\bar{\theta}C_{i}-\log\left(\theta^{*}\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\right\},$
$\bar{\beta}^{(t)}_{i}(\xi;\eta)=\exp\left\{-\mathrm{CI}t-\frac{1}{2}\log t+(1-\theta^{*})\eta+\bar{\theta}C_{i}-\log\left((1-\theta^{*})\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\right\}.$
@@
$\mathsf{P}_{\mathsf{e},i}^{(t)*}(\xi)\leq$ $\;\xi_{0}\alpha^{(t)}_{i}(\xi;\eta)+\xi_{1}\beta^{(t)}_{i}(\xi;\eta)\leq\xi_{0}\bar{\alpha}^{(t)}_{i}(\xi;\eta)+\xi_{1}\bar{\beta}^{(t)}_{i}(\xi;\eta)$
@@
$=$ $\;\xi_{0}^{1-\theta^{*}}\xi_{1}^{\theta^{*}}\exp\left\{-\mathrm{CI}t-\frac{1}{2}\log t+\bar{\theta}C_{i}-\log\left(\theta^{*}(1-\theta^{*})\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\right\}$
@@
A-F Proof of Lemma 2
The proof is similar to the one for Proposition 3 in [14]. We have
$\left(\sum_{j=1}^{n}\left|\left[W^{\tau}\right]_{ij}-\pi_{j}\right|r_{j}\right)^{2}\leq\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left|\left[W^{\tau}\right]_{ij}-\pi_{j}\right|^{2}\right)\left(\sum_{j=1}^{n}\pi_{j}r_{j}^{2}\right)$ (Cauchy-Schwarz)
and
$\;\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left|\left[W^{\tau}\right]_{ij}-\pi_{j}\right|^{2}=\sum_{j=1}^{n}\left(\frac{1}{\pi_{j}}\left(\left[W^{\tau}\right]_{ij}\right)^{2}-2\left[W^{\tau}\right]_{ij}+\pi_{j}\right)$
Following the remaining part in the proof for Proposition 3 in [14], Lemma 2 is proved.
A-G Gaussian Case
@@
$\mathcal{H}_{0}:$ $\;\;X^{(t)}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(-\mu,\sigma^{2}),$
$\mathcal{H}_{1}:$ $\;\;X^{(t)}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\mu,\sigma^{2}),$
@@
$\mathcal{H}_{0}:$ $\;\;\mathcal{N}(-\tilde{\mu},\tilde{\sigma}^{2}),$
$\mathcal{H}_{1}:$ $\;\;\mathcal{N}(\tilde{\mu},\tilde{\sigma}^{2})$
@@
$=$ $\;\frac{1}{A\sqrt{2\pi}}\exp\left\{\frac{-2\mu}{\sigma}\underbrace{\frac{\left(\sum_{\tau,j=1}^{t,n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}}{\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}}}_{B}+\frac{2\mu}{\sigma}\underbrace{\frac{\left(\sum_{\tau,j=1}^{t,n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)}{\sqrt{\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}}}}_{C}\lambda-\frac{\lambda^{2}}{2}\right\}.$ (35)
@@
$=$ $\;\frac{tn+2\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)+\frac{1}{tn}\left[\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)\right]^{2}}{1+\frac{1}{tn}\sum_{\tau,j=1}^{t,n}\left[\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}-1\right]}$
$=$ $\;\left[tn+2\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)+o\left(\frac{1}{t}\right)\right]\left[1-\frac{1}{tn}\sum_{\tau,j=1}^{t,n}\left[\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}-1\right]+o\left(\frac{1}{t}\right)\right]$
$=$ $\;\left[tn+2\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)\right]\left[1-\frac{2}{tn}\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)-\frac{1}{tn}\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)^{2}\right]+o(1)$
@@
$B\geq$ $\;tn-\sum_{\tau=1}^{t}\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}\right)\rho^{2\tau}+o(1)=tn-\left(\frac{\rho^{2}}{1-\rho^{2}}\right)\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}\right)+o(1).$
@@
$C=B^{\frac{1}{2}}=\sqrt{tn-\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)^{2}+o(1)}=\sqrt{tn}+o(1).$
@@
$A=\left(\frac{2\mu\sqrt{n}}{\sigma}\sqrt{t}-\lambda+o(1)\right)=\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{t}}(1+o(1))=\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{t}}e^{o(1)}.$
@@
$\beta^{(t)*}_{i}(r^{*},\epsilon)\leq\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{2\pi t}}\exp\left\{-\frac{2\mu n}{\sigma}t+\frac{2\mu\lambda\sqrt{n}}{\sigma}\sqrt{t}-\frac{\lambda^{2}}{2}+\frac{2\mu}{\sigma}\left(\frac{\rho^{2}}{1-\rho^{2}}\right)\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}\right)+o(1)\right\}.$
@@
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)=\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{2\pi t}}\exp\left\{-\frac{2\mu n}{\sigma}t+\frac{2\mu\lambda\sqrt{n}}{\sigma}\sqrt{t}-\frac{\lambda^{2}}{2}+o(1)\right\}.$

### v2

@@ A-E Proof of Theorem 5
For the product distribution of the nodes, first let $x$ denote $(x^{(1)}_{1},x{(t)}_{2},\dots,x^{(t)}_{n})\in\mathcal{X}^{t}$, and we have
@@
$\beta^{(t)}_{i}(\xi;\eta)$
$=P_{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(X^{(t)}_{j})}{P_{j,0}(X^{(t)}_{j})}\leq\eta\right\}$
@@
[…]
$\quad\cdot\mathbbm{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}\leq\eta\right\}.$ (34)
@@
$\sum_{j=1}^{n}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,\theta}(x^{(t)}_{j})}$
$=(1-\theta^{*})\left(\sum_{j=1}^{n}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}\right)+\mathrm{CI}(P_{0},P_{1}).$
@@
[…]
$=\sum_{x\in\mathcal{X}^{t}}\left\{\begin{array}[]{l}P^{\otimes t}_{\theta^{*}}(x)e^{(1-\theta^{*})\left(\sum_{\tau=1}^{t}\sum_{j=1}^{n}\log\frac{P_{j,0}(x^{(t)}_{j})}{P_{j,\theta}(x^{(t)}_{j})}\right)-\mathrm{CI}t}\\ \cdot\mathbbm{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}\leq\eta\right\}\end{array}\right\}.$ (37)
Similar to the trick in Appendix A-D, we have
@@
$=\sum_{\tau=1}^{t}\sum_{j=1}^{n}(1-\theta^{*})\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}$
$\quad+\sum_{\tau=1}^{t}\sum_{j=1}^{n}\left(1-(1-\theta^{*})\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}$
$\leq\sum_{\tau=1}^{t}\sum_{j=1}^{n}(1-\theta^{*})\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x^{(t)}_{j})}{P_{j,0}(x^{(t)}_{j})}$
$\quad+\underbrace{(1-\theta^{*})\frac{\rho}{1-\rho}\sqrt{\frac{1-\pi_{i}}{\pi_{i}}\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}L^{2}_{j}\right)}}_{C_{i}}.$
@@
$=e^{-\mathrm{CI}t+(1-\theta^{*})C_{i}}\!\sum_{x\in\mathcal{X}^{t}}P^{\otimes t}_{\theta^{*}}(x)e^{(1-\theta^{*})w(x)}\mathbbm{1}\left\{w(x)\leq\eta\right\}.$ (38)
@@
By Lemma 3, we know that $Y^{(t)}_{i}$’s CDF converges to the one of the standard normal distribution with some remaining terms, and so far we have equation (38) become
$e^{-\mathrm{CI}t+(1-\theta^{*})C_{i}}$
$\quad\cdot\sum_{y\in\mathcal{Y}}\Bigg{(}\begin{array}[]{l}P_{Y^{(t)}_{i}}(y)e^{(1-\theta^{*})S^{(t)}_{n}\sqrt{t}y}\\ \mathbbm{1}\left\{(1-\theta^{*})S^{(t)}_{n}\sqrt{t}yx\leq(1-\theta^{*})\eta\right\}\end{array}\Bigg{)}$
$=e^{-\mathrm{CI}t+(1-\theta^{*})\eta+(1-\theta^{*})C_{i}}\sum_{z\leq 0}P_{Y^{(t)}_{i}}\left(\frac{z+(1-\theta^{*})\eta}{(1-\theta^{*})S^{(t)}_{n}\sqrt{t}}\right)e^{z}.$
@@
using the similar method in Appendix A-D. Thus, the type-II error probability is
$\beta^{(t)}_{i}(\xi;\eta)$
$=\exp\left\{\begin{array}[]{l}-\mathrm{CI}t-\frac{1}{2}\log t+(1-\theta^{*})\eta+(1-\theta^{*})C_{i}\\ -\log\left((1-\theta^{*})\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\end{array}\right\}.$ (41)
@@
$\alpha^{(t)}_{i}(\xi;\eta)$
$=\exp\left\{\begin{array}[]{l}-\mathrm{CI}t-\frac{1}{2}\log t-\theta^{*}\eta+\theta^{*}C_{i}\\ -\log\left(\theta^{*}\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\end{array}\right\}.$ (44)
@@
$\bar{\alpha}^{(t)}_{i}(\xi;\eta)$ $=\exp\left\{\begin{array}[]{l}-\mathrm{CI}t-\frac{1}{2}\log t-\theta^{*}\eta+\bar{\theta}C_{i}\\ -\log\left(\theta^{*}\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\end{array}\right\},$
$\bar{\beta}^{(t)}_{i}(\xi;\eta)$ $=\exp\left\{\begin{array}[]{l}-\mathrm{CI}t-\frac{1}{2}\log t+(1-\theta^{*})\eta+\bar{\theta}C_{i}\\ -\log\left((1-\theta^{*})\sigma\right)-\frac{1}{2}\log(2\pi)+o(1)\end{array}\right\}.$
@@
$\mathsf{P}_{\mathsf{e},i}^{(t)*}(\xi)$
$\leq\xi_{0}\alpha^{(t)}_{i}(\xi;\eta)+\xi_{1}\beta^{(t)}_{i}(\xi;\eta)$
$\leq\xi_{0}\bar{\alpha}^{(t)}_{i}(\xi;\eta)+\xi_{1}\bar{\beta}^{(t)}_{i}(\xi;\eta)$
@@
$=\xi_{0}^{1-\theta^{*}}\xi_{1}^{\theta^{*}}\exp\left\{\begin{array}[]{l}-\mathrm{CI}t-\frac{1}{2}\log t+\bar{\theta}C_{i}\\ -\log\left(\theta^{*}(1-\theta^{*})\sigma\right)-\frac{1}{2}\log(2\pi)\\ +o(1)\end{array}\right\}$
@@
A-F Gaussian Case
@@
[…]
@@
$\mathcal{H}_{0}:$ $\;\;\mathrm{Normal}(-\tilde{\mu},\tilde{\sigma}^{2}),$
$\mathcal{H}_{1}:$ $\;\;\mathrm{Normal}(\tilde{\mu},\tilde{\sigma}^{2})$
@@
$=\frac{1}{A\sqrt{2\pi}}\exp\left\{\frac{-2\mu}{\sigma}B+\frac{2\mu}{\sigma}C\lambda-\frac{\lambda^{2}}{2}\right\},$ (46)
where the shorthand notations
$B=\frac{\left(\sum_{\tau,j=1}^{t,n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}}{\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}},\ C=\frac{\left(\sum_{\tau,j=1}^{t,n}\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)}{\sqrt{\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}}}.$
@@
$=\frac{\left(\begin{array}[]{l}tn+2\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)\\ +\frac{1}{tn}\left[\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)\right]^{2}\end{array}\right)}{1+\frac{1}{tn}\sum_{\tau,j=1}^{t,n}\left[\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}-1\right]}$
$=\left[tn+2\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)+o\left(\frac{1}{t}\right)\right]$
$\quad\cdot\left[1-\frac{1}{tn}\sum_{\tau,j=1}^{t,n}\left[\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}\right)^{2}-1\right]+o\left(\frac{1}{t}\right)\right]$
$=\left[tn+2\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)\right]$
$\quad\cdot\left[\begin{array}[]{l}1-\frac{2}{tn}\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)\\ -\frac{1}{tn}\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)^{2}\end{array}\right]+o(1)$
@@
$B$ $\geq tn-\sum_{\tau=1}^{t}\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}\right)\rho^{2\tau}+o(1)$
$=tn-\left(\frac{\rho^{2}}{1-\rho^{2}}\right)\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}\right)+o(1).$
@@
$C=B^{\frac{1}{2}}$ $=\sqrt{tn-\sum_{\tau,j=1}^{t,n}\left(\frac{[W^{\tau}]_{ij}}{\pi_{j}}-1\right)^{2}+o(1)}$
$=\sqrt{tn}+o(1).$
@@
$A$ $=\left(\frac{2\mu\sqrt{n}}{\sigma}\sqrt{t}-\lambda+o(1)\right)$
$=\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{t}}(1+o(1))=\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{t}}e^{o(1)}.$
@@
$\beta^{(t)*}_{i}(r^{*},\epsilon)$
$\leq\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{2\pi t}}\exp\left\{\begin{array}[]{l}-\frac{2\mu n}{\sigma}t+\frac{2\mu\lambda\sqrt{n}}{\sigma}\sqrt{t}-\frac{\lambda^{2}}{2}\\ +\frac{2\mu}{\sigma}\big{(}\frac{\rho^{2}}{1-\rho^{2}}\big{)}\big{(}\frac{\pi_{i}}{1-\pi_{i}}\big{)}\!\sum\limits_{j=1}^{n}\frac{1}{\pi_{j}}\\ +o(1)\end{array}\right\}.$
@@
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)$
$=\frac{1}{\frac{2\mu\sqrt{n}}{\sigma}\sqrt{2\pi t}}\exp\left\{-\frac{2\mu n}{\sigma}t+\frac{2\mu\lambda\sqrt{n}}{\sigma}\sqrt{t}-\frac{\lambda^{2}}{2}+o(1)\right\}.$